Circular No. 6810

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1997ff AND 1997fg R. L. Gilliland, Space Telescope Science Institute; and M. M. Phillips,

Cerro Tololo Interamerican Observatory, report their discovery of two apparent supernovae in the Hubble Deep Field (HDF). The objects were detected in second-epoch exposures of the HDF obtained with WFPC2 on the Hubble Space Telescope (+ F814W filter) between 1997 Dec. 23.6 and 27.0 UT; they were not present in images taken on 1995 Dec. 25. SN 1997ff is located 0′′.11 west and 0′′.11 south of the center of a galaxy designated by Williams et al. (1996, A.J. 112, 1335) as 4-403.0, whose position is given as α = 12^{h}36^{m}44^{s}.11, δ = +62^{o}12′44′′.8 (equinox 2000.0). SN 1997fg is located 0′′.31 east and 0′′.31 south of the center of 3-221.0, which has position end figures 57^{s}.67, 13′15′′.3 (and redshift 0.952). Magnitudes were obtained by breaking the new series of HDF observations into three sets (each of total exposure 21 000 s) and were derived with reference to the I-band magnitudes of stars in the HDF from Flynn et al. (1996, Ap.J. 466, L55): SN 1997ff, Dec. 23.83 UT, I = 26.77 ± 0.14; 25.78, 27.20 ± 0.21; 26.74, 26.85 ± 0.15. SN 1997fg, Dec. 23.83, I = 25.93 ± 0.07; 25.78, 26.04 ± 0.07; 26.74, 26.02 ± 0.07. SN 1997ff is a 9σ detection, while SN 1997fg is > 20σ.

NGC 2060 F. E. Marshall, Goddard Space Flight Center (GSFC); J. Middleditch,

Los Alamos National Laboratory; W. Zhang, GSFC; and E. V. Gotthelf, GSFC and Universities Space Research Association, report the discovery of pulsations from NGC 2060 = N157B, a Crab-like supernova remnant in the Large Magellanic Cloud (cf. Henize 1956, Ap.J. Suppl. 2, 331; Mathewson et al. 1983, ibid. 51, 345): “An observation with the Rossi X-ray Timing Explorer on 1996 Oct. 12 revealed pulsations in the band 2–10 keV with a period of 16.114712 ± 0.000003 ms. Pulsations were also seen in the band 10–25 keV. Analysis of archival data acquired with the ASCA Xray Observatory confirms the RXTE detection and the identification with N157B. Using a subsequent RXTE detection of the pulsar on 1996 Dec. 22, and ASCA detections on 1993 June 13, Aug. 29, Sept. 23, and 1995 Nov. 6, we derive a period derivative of (5.124 ± 0.003) × 10^{−14} s/s. This appears to be the most rapidly rotating pulsar yet seen that has not been spun up since its birth. The pulsar is likely associated with the compact x-ray source in N157B (Wang and Gotthelf 1998, Ap.J., in press; preprint available at

http://xxx.lanl.gov/abs/astro-ph/?9708087). Observations at other

wavelengths are encouraged.”

1998 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green